OpenStudy (anonymous):

54 caitlynmcdaniel Medals 0 3.) a jet took one hour longer flying to Lincoln from Adams at 800 km/h than to return at 1200 km/h find the distance from Lincoln to Adams

OpenStudy (radar):

Regardless of the distance, we know that whatever it is, it is the same both ways. You agree?

OpenStudy (anonymous):

agree

OpenStudy (anonymous):

the teacher wants a chart then equations. rate fist two boxes down time second two boxes down next distance second two boxes down... r times time=distance

OpenStudy (anonymous):

u there?????????????????? @radar

OpenStudy (radar):

You will have to make your own chart.

OpenStudy (anonymous):

to clearify the distance is 2400 km that's the answer?

OpenStudy (radar):

Lets call that distance D. Also lets call the time it takes to fly from Lincoln to Adams. T The time then to fly to Lincoln from Adams is T + 1. This is all information provided by the problem. In addition we are given speed each way. Distance is equal to speed times the time. So lets set up our equations: D = 800(T + 1) (Lincoln from Adams) and D = 1200T (Adams *from Lincoln) D = D so we can now write 800(T+1) = 1200T Now the algebra. 800T + 800 = 1200T 800T - 1200T = -800 -400T = -800 T = 2 hrs. T + 1 = 3 (hrs Now the distance is a piece of cake. D=800(2 + 2) =800 (3) =2400 km and D the other way would be 1200T=1200 (2) = 2400 km You see it is the same distance going each direction.

OpenStudy (radar):

Yes, both ways lol.

OpenStudy (anonymous):

thanks I wrote every step down.. will u help with 4,5,9??????

OpenStudy (radar):

What kinds of animal is in the picture.

OpenStudy (anonymous):

what????????????

OpenStudy (radar):

Your logo, or icon ??

OpenStudy (radar):

Is it a rabbit?

OpenStudy (anonymous):

will u help with numbers 4,5,9??? ohh my baby chihuaha (: hes in my room with me now

OpenStudy (radar):

I am sorry but I have to work on a form for the court. I will be back in a few hours.

OpenStudy (anonymous):

okie dokie